# HiFi-Stream: Streaming Speech Enhancement with Generative Adversarial Networks

### v1

@@ Abstract
Speech Enhancement techniques have become core technologies in mobile devices and voice software simplifying downstream speech tasks. Still, modern Deep Learning (DL) solutions often require high amount of computational resources what makes their usage on low-resource devices challenging. We present HiFi-Stream, an optimized version of recently published HiFi++ model. Our experiments demonstrate that HiFiStream saves most of the qualities of the original model despite its size and computational complexity: the lightest version has only around 490k parameters which is 3.5x reduction in comparison to the original HiFi++ making it one of the smallest and fastest models available. The model is evaluated in streaming setting where it demonstrates its superior performance in comparison to modern baselines.
@@ I Introduction
Speech enhancement (SE) techniques are generally used as a preprocessing tool for speech downstream tasks such as speech recognition. The main goal of speech enhancement is to remove noise and reverberations coming from the actual environment where the audio is recorded. Mobile and wearable smart devices such as smartphones and smartwatches naturally face strongly corrupted speech as an audio input. Studies [1], [2] show that the quality of automatic speech recognition (ASR) solutions benefits from using an additional speech enhancement model making it an important part of speech processing pipeline. In the area of SE, the state-of-the-art Deep Learning (DL) approaches [3], [4], [5] demonstrated that DL can manage very hard audio conditions in speech enhancement and source separation problems. Nonetheless, the line of research in these models was mainly achieving quality at all costs resulting in dozens of millions parameters. On the other hand, such models require powerful computational resources which are unlikely to be placed in a mobile or wearable device. To get the scale, DEMUCS [3] hit the relevant realtime constraints as a result of significant optimization efforts. The architecture has around 60 M parameters and requires 38.1 GMACs (Multiply–ACcumulate operations) per second of audio in 16 kHz [6], [7]. The problem of constructing architectures for various speech processing tasks available on low-resource devices has become the focus of research efforts in recent years [8].
Generative Adversarial Networks (GANs) quickly showed their superior quality in text-to-speech problems [9], but still the right-away application of GANs for SE problems [10],[11],[12] resulted in very computationally expensive models. Recently [6] it was shown that the number of parameters used in GANs for SE can be reduced more than 10 times suggesting that with more careful design there is an opportunity to propose even more efficient solution with almost the same performance on the benchmarks.
In our work we propose HiFi-Stream, an optimized version of HiFi++ [6] introducing structural and technical modifications. The experimental set up, pipeline and model configurations are given in the github repository: https://github.com/KVDmitrieva/source_sep_hifi.
@@
    We show that the model can be reduced almost 3.5 times in parameters and around 30% in MACs making it one of the available quality-complexity trade-offs.
@@
    We also investigate the performance of HiFiStream in an audio stream simulation where the audio is given in chunks. HiFi-Stream demonstrates superior quality to its direct competitors : HiFi++[6] and FFC-SE[7]; both in performance and computational complexity without additional streaming training.
Figure 1: HiFi++ and Hifi-Stream architectures: the new model reduced the number of parameters, got new MaskNet block, redesigned HiFi upsampler and removed SpectralUNet
@@ II Related Work
Speech enhancement and speech separation methods generally can be classified into two branches of research. Time-frequency (TF) methods are based on the usage of the spectrograms produced by the Short-Time Fourier Transform (STFT). Models take a spectrogram as input and outputs the cleaned spectrogram which is transformed into a cleaned waveform via inverse STFT (iSTFT). The TF methods are represented in particular by [14, 4, 1].Time-domain (TD) methods got their popularity just in recent years [3, 15, 16, 5, 17]. Most of these models by construction require millions of parameters, the most efficient of these are DPRNN models with around 2-6 M parameters [15]. For these reason in our research we focus on TF-methods.
Streaming processing presents additional constraints [8]. The first is computational complexity reflected by the number of parameters and processing time. The second is causality: the model is restricted to use only limited context. So, the researchers are focused on constructing streaming-aware versions of state-of-the-art solutions some of which include ReSepformer [18] and SkiM [19]. It was recently shown that regression transformer-based solutions can be implemented in low-resource setting but with considerable efforts invested into the attention mechanisms [8].
The generative look at the problem presented by SEGAN [10] later developed into HiFi-GAN solutions [11] and [12]. It was suggested to train and design the model as GAN where the generator predicts the clean waveform. This was reviewed in HiFi++ [6] with more structured design. The model achieved the baseline results despite significantly reduced model complexity. Several other solutions of the same scale were proposed later [20, 7, 21]. In [6], [7] the main goal was to reduce the model size but computational complexity was not addressed. Our aim is to improve also the computational complexity of the SE, and we improved the results of [6, 7] in this regard, see Fig. 3.
@@ III Methodology
The problem of speech enhancement is to get an estimate $\hat{x}\in\mathbb{R}^{N_{ch}\times T}$ of the clean speech signal $x\in\mathbb{R}^{N_{ch}\times T}$ with $N_{ch}$ channels from its corrupted representation
$y=f\star x+\varepsilon,$
where $f\in\mathbb{R}^{N_{ch}\times k}$ represents an impulse response kernel (the effects of the environment acoustics, reverberation and recording device) which is convolved with the clean signal and $\varepsilon\in\mathbb{R}^{N_{ch}\times T}$ is an additive noise component. In our work we consider a monaural problem, i.e. the signal has only one channel. Originally GAN approaches mostly [10], [11], [12] used time-domain approach requiring direct signal reconstruction. It turned out that such approaches typically have large number of parameters: one needs a sophisticated enough encoder and decoder to obtain a representative set of features from raw audio. So, HiFi++ [6] and our work in this aspect take a route of TF-methods rising from original vocoding applications [9]. Classic TF approaches usually propose to estimate a magnitude mask $M(y)\in\mathbb{R}^{F\times M}$ with values from 00 to $1$ for signal spectrogram $s=STFT(y)\in\mathbb{C}^{F\times M}$ and then estimate the clean signal as
$\hat{x}=iSTFT(M(y)\odot s).$
In such a way, the phase component is bypassed to have a well-defined inverse STFT. On a high level, in HiFi++ the signal is first re-synthesized in several audio channels and then is processed by the masking pipeline producing the spectral mask.
The generator depicted on the Fig. 1 consists of several blocks: STFT-Mel transform, Spectral U-Net, HiFi, Wave U-Net and SpectralMaskNet. The idea is to train the generator to predict clean waveform and discriminator decides whether the audio is clean. We used the same loss construction and discriminators as in [6, 9] and focus our main efforts on the generator structure.
@@
HiFi block is the largest core block in the model. When trained, it transforms mel-spectrogram $m=Mel(s)$ of the signal $x$ into multichannel audio $x^{\prime}\in\mathbb{R}^{n_{ch}\times T}$. The channels can be interpreted as signal and noise layers very similar to those observed in the output of source separation [15]. We experimented with usage of 2D convolutions instead of 1D and varying parameters, which gave a drastic reduction in almost -400k parameters but in the same time GMACs became slightly higher with 2D convolutions. Still, the usage of 2D convolutions even with smaller kernel size increased the performance, since the model better exploits the spatio-temporal correlations in the spectrogram.
Wave UNet
The block is the main post-processing unit responsible for fixing the signal in time domain [6] by removing robo-voice artifacts typical for vocoding [9]. With the help of input waveform and multichannel vocoder output it effectively fixes the phase shifts. The Wave UNet design is taken from [6].
@@
SpectralMaskNet is another key component in the model since it generates the spectral magnitude mask $M(y)$. The original HiFi++ used a U-Net processing unit which can be improved both in terms of parameters and computational complexity. The mask of gains is essentialy a tensor of the same size as spectrogram where larger values depict higher amplitude. Such specification very natural for attention architectures [22] but it is known that despite attention design allows more expressive models, it is computationally expensive for streaming [8]. We came up with a cheaper attention alternative from RawNet2 [23] called Feature Map Scaling (FMS), known also as Squeeze-Excitation [24]. Instead of original UNet we used FMS Blocks with affine aggregation (see Fig. 2) in a ResNet structure. FMS acts as attention mechanism with reduced complexity. The usage of FMS ResNet gave 20x reduction in the MaskNet block.
Figure 2: FMS Block inspired by [23] takes the input tensor $C\in\mathbb{R}^{F\times T}$, computes the vector of scaling coefficients $S\in\mathbb{R}^{T}$ and applies them filter-wise and computes the result via affine aggregation.
@@
Model WV-MOS SI-SDR SDR STOI PESQ CSIG CBAK COVL nParams GMACs/sec
Ground Truth 4.500 - - 1.000 4.640 5 5 5 - -
Input 2.990 8.440 8.54 0.790 1.970 3.340 2.820 2.740 - -
SEGAN[10] - - - - 2.160 3.480 2.940 2.80 - -
MetricGAN+[25] 3.913 8.162 14.734 0.853 3.130 3.971 3.123 3.549 2.700M 28.500
DEMUCS(non-caus.)[3] 4.453 18.585 19.603 0.893 3.015 4.348 3.597 3.715 30.100M 38.1
HiFiGAN-SE* [11] - - - - 2.940 4.070 3.070 3.490 - -
HiFiGAN-2* [12] - - - - 3.110 4.370 3.540 3.740 - -
FFC-AE-V0 [7] 4.339 16.313 17.291 0.865 2.669 4.061 3.296 3.373 0.422M 4.064
FFC-AE-V1 [7] 4.402 16.899 17.484 0.893 2.868 4.238 3.497 3.573 1.665M 15.672
FFC-UNet [7] 4.421 17.851 18.702 0.894 2.962 4.349 3.510 3.682 7.673M 35.552
HiFi++ [6] 4.385 17.954 18.936 0.890 2.906 4.271 3.483 3.609 1.706M 2.746
HiFi++(reprod.) 4.314 13.400 15.661 0.866 2.595 3.998 3.173 3.305 1.617M 2.096
HiFi w/o spec 4.278 12.808 16.155 0.873 2.703 4.094 3.166 3.413 1.356M 2.007
HiFi-2dMRF 4.130 13.455 15.812 0.865 2.466 3.865 3.075 3.172 0.677M 2.084
HiFi-fms (HiFi-Stream) 4.320 14.067 15.924 0.865 2.701 4.031 3.220 3.372 1.174M 1.895
HiFi-fms+2dMRF (HiFi-Stream2D) 4.312 14.762 17.213 0.864 2.652 4.027 3.167 3.345 0.497M 1.973
@@ IV Experimental Setup
Model WV-MOS SI-SDR SDR STOI PESQ CSIG CBAK COVL
Ground Truth 4.500 - - 1.000 4.640 5 5 5
Input 2.990 8.440 8.54 0.790 1.970 3.340 2.820 2.740
DEMUCS(caus.dns)[3] 4.600 17.804 19.310 0.881 2.657 3.823 3.231 3.236
FFC-AE-V0 [7] 4.209 12.664 13.820 0.789 1.996 3.502 2.807 2.742
FFC-AE-V1 [7] 4.365 15.147 15.806 0.860 2.424 3.906 3.214 3.172
FFC-UNet [7] 4.378 15.683 16.467 0.861 2.452 3.974 3.225 3.222
HiFi++ [6] 4.263 14.669 15.264 0.930 2.365 3.838 3.162 3.103
HiFi++(reprod.) 4.253 13.014 15.383 0.841 2.263 3.715 2.917 2.992
HiFi w/o spec 4.059 11.839 14.746 0.791 1.932 3.200 2.671 2.559
HiFi-2dMRF 3.995 12.771 14.861 0.806 1.931 3.259 2.726 2.587
HiFi-fms (HiFi-Stream) 4.247 13.328 15.123 0.845 2.295 3.667 2.860 2.981
HiFi-fms+2dMRF (HiFi-Stream2D) 4.296 14.189 16.328 0.847 2.286 3.680 2.925 2.982
Figure 3: Performance metric (COVL) against the computational complexity and number of parameters for compared models (streaming evaluation)
IV-A Dataset
We used Valentini VCTK [13] dataset for experiments. The dataset is a selection of 30 speakers from the VoiceBank corpus with 28 and 2 speakers set for the training and test respectively. The original sample rate is 48 kHz, we resampled the audio to 16 kHz for comparison with other solutions. All audios were cut to $3$ seconds as in [6].
@@ IV-A Dataset
For evaluation we used perceptual evaluation metrics: CSIG, CBAK, COVL, WB-PESQ, STOI, WV-MOS [6]. Also, we present objective evaluation metrics: SDR and SI-SDR. The key aspect of our research is model complexity, so we also demonstrate the complexity metrics: number of parameters and GMACs per second of audio measured with pytorch profiling library thop.
@@ IV-B Training details
All models were trained for 150 epochs with learning rate 2e-4 on the given dataset with batch size 32 using AdamW optimizer (see the repository for the exact details). For HiFi++ we trained our own model (marked as reprod.) with the same computational budget but also added the metrics for the checkpoint given by [6].

### v2

@@ Abstract
Speech Enhancement techniques have become core technologies in mobile devices and voice software. Still, modern deep learning solutions often require high amount of computational resources what makes their usage on low-resource devices challenging. We present HiFi-Stream, an optimized version of recently published HiFi++ model. Our experiments demonstrate that HiFi-Stream saves most of the qualities of the original model despite its size and computational complexity improved in comparison to the original HiFi++ making it one of the smallest and fastest models available. The model is evaluated in streaming setting where it demonstrates its superior performance in comparison to modern baselines.
@@ I Introduction
Speech enhancement (SE) techniques are generally used as a preprocessing tool for speech downstream tasks such as speech recognition [12, 21]. The main goal of SE is to remove noise and reverberations coming from the actual environment where the audio is recorded. In the area of SE, the state-of-the-art Deep Learning (DL) approaches demonstrated that DL can manage very hard audio conditions in SE and source separation problems. Typically DL models require powerful computational units which are unlikely to be placed in a mobile or wearable device. Construction of architectures for speech processing tasks better suited for low-resource applications has recently become the focus of research efforts [23].
In our work we propose HiFi-Stream models, two optimized versions of HiFi++ [1] introducing structural and technical modifications. The experimental set-up, pipeline and model configurations are given in the github repository: https://github.com/KVDmitrieva/source_sep_hifi.
@@
    HiFi-Stream models are reduced at least nearly 40% in parameters and around 10% in MACs making them a viable quality-complexity trade-offs.
@@
    We evaluate the trained HiFi-Stream models in audio stream simulation where the audio is processed in chunks. HiFi-Stream outperforms its direct competitors: HiFi++[1] and FFC-SE[15]; without additional streaming training. We found the new attention-like module for mask construction to be the key ingredient resulting in model robustness in limited context.
Figure 1: HiFi++ and Hifi-Stream architectures: the new model reduced the number of parameters, got new MaskNet block, redesigned HiFi upsampler and removed Spectral U-Net.
@@ II Related Work
SE and speech separation methods generally could be classified into time-frequency (TF) [2, 6, 12] and time-domain (TD) [3, 10, 20] models. Despite the growing interest in time-domain models, they remain to be very complex with at least millions of parameters [10]. Studies demonstrate that TF methods have advantage over TD methods because magnitude spectrogram serves as very expressive and in the same time is processed very efficiently [2, 6].
Streaming processing presents several specific constraints [23]. The first is computational complexity reflected by the number of parameters and processing time. The second is causality: the model is restricted to use only limited context. So, the researchers are focused on constructing streaming-aware versions of state-of-the-art solutions based on transformer architectures [8]. It was recently shown that transformer-based solutions can be implemented in low-resource setting but with considerable efforts invested into the attention mechanisms [23].
The generative look at the problem presented by SEGAN [11] later developed into HiFi-GAN solutions [16, 17]. It was suggested to train and design the model as GAN where the generator predicts the clean waveform. This was reviewed in HiFi++ [1] with more structured design. The model achieved the baseline results despite significantly reduced model complexity. Several other solutions of the same scale were proposed later [9, 15, 24]. In [1], [15] the main goal was to reduce the model size but computational complexity was not addressed. Our aim is to improve also the computational complexity of the SE, and we improved the results of [1, 15] in this regard, see Fig. 3.
@@ III Methodology
The problem of speech enhancement is to get an estimate of the clean speech signal $x\in\mathbb{R}^{N_{ch}\times T}$ with $N_{ch}$ channels from its corrupted representation
$y=f\star x+\varepsilon,$ (1)
where $f\in\mathbb{R}^{N_{ch}\times k}$ represents an impulse response (IR) kernel (the effects of the environment acoustics, reverberation and recording device) and $\star$ depicts a convolution operation. In such notation, the IR kernel is convolved with the clean signal and $\varepsilon\in\mathbb{R}^{N_{ch}\times T}$ serves as an additive noise component. In our work we consider a problem with $N_{ch}=1$ though the modification for other cases can be done straightforwardly. Originally GAN approaches mostly used time-domain approach requiring direct signal reconstruction and implying large number of parameters since one needs quite complex encoder and decoder to obtain a representative set of features from raw audio. So, we in our work take a route of TF-methods rising from original vocoding applications. Classic TF approaches usually propose to estimate a magnitude mask $M(y)\in\mathbb{R}^{F\times M}$ with values from 00 to $1$ for signal spectrogram $s=STFT(y)\in\mathbb{C}^{F\times M}$ and then estimate the clean signal as
$\hat{x}=iSTFT(M(y)\odot s)$ (2)
with $\odot$ denoting coordinate-wise product. In such a way, the phase component is bypassed to have a well-defined inverse STFT. At a high level, in HiFi++, the signal is first re-synthesized across several audio channels and then processed by the masking pipeline, which produces the spectral mask.
The generator depicted on the Fig. 1 consists of several blocks: STFT-Mel transform, Spectral U-Net, HiFi, Wave U-Net and SpectralMaskNet. The idea is to train the generator to predict clean waveform while discriminator decides whether the audio is clean. The model operates in steps.
@@
HiFi block is the largest core block in the model responsible for upsampling the magnitude spectrogram back to signal domain. When trained, it transforms mel-spectrogram $m=Mel(s)$ of the signal $x$ into multichannel audio $x^{\prime}\in\mathbb{R}^{n_{ch}\times T}$. The channels can be interpreted as signal and noise layers very similar to those observed in the output of source separation. We experimented with using 2D convolutions across time and frequency dimensions instead of 1D (across frequency only) and varying kernel size and channel parameters, which resulted in a drastic reduction of nearly 400k parameters since the number of out channels in convolutions reduced around 4 to 8 times. However, at the same time, GMACs slightly increased because of higher computational complexity of 2D convolutions. Still, we increased the performance, since the model better exploits the frequency-temporal correlations in the spectrogram.
Wave U-Net
is the main post-processing unit responsible for fixing the signal in time domain by removing robo-voice artifacts typical for vocoding [7]. With the help of input waveform concatenated with multichannel HiFi output it effectively fixes the phase shifts. The Wave U-Net design is taken from [1].
@@
SpectralMaskNet is another key component in the model since it generates the spectral magnitude mask $M(y)$. HiFi++ used a U-Net convolutional processing unit. It outputs mask of gains, essentialy a tensor of the same size as spectrogram. Such processing is very natural for attention architectures. However, despite allowing for more expressive models, attention architecture is computationally expensive for streaming [23]. We came up with a cheaper attention alternative from RawNet2 [22] called Feature Map Scaling (FMS), known also as Squeeze-Excitation [5]. For masking we constructed FMS Blocks with affine aggregation (see Fig. 2) in a ResNet structure. FMS acts as more complexity-friendly attention alternative. This choice resulted in 20x reduction in MaskNet block in comparison to the original U-Net because of less convolution channels. In new MaskNet with just $(8,16,8)$ channels with FMS transform in between we achieved better computational efficiency due to the attention nature of FMS and absence of the bottleneck. Our key finding is that attention-like mechanism showed its robustness in the streaming setting with reduced context (see Table II).
Figure 2: FMS Block inspired by [22] takes the input tensor $C\in\mathbb{R}^{F\times T}$, computes the vector of scaling coefficients $S\in\mathbb{R}^{T}$ and applies them across filters and computes the result tensor $C^{\prime}_{ft}=C_{ft}S_{t}+C_{ft}$. This method turns out to be a key component of model robustness in streaming mode and serves as temporal attention.
@@
Model SI-SDR (dB) $\uparrow$ SDR (dB) $\uparrow$ STOI $\uparrow$ PESQ $\uparrow$ CSIG $\uparrow$ CBAK $\uparrow$ COVL $\uparrow$ nParams GMACs/sec
Ground Truth - - 1.000 4.640 5 5 5 - -
Input 8.440 8.54 0.790 1.970 3.340 2.820 2.740 - -
SEGAN[11] - - - 2.160 3.480 2.940 2.80 - -
MetricGAN+[4] 8.162 14.734 0.853 3.130 3.971 3.123 3.549 2.700M 28.500
DEMUCS(non-caus.)[3] 18.585 19.603 0.893 3.015 4.348 3.597 3.715 30.100M 38.1
HiFiGAN-SE* [16] - - - 2.940 4.070 3.070 3.490 - -
HiFiGAN-2* [17] - - - 3.110 4.370 3.540 3.740 - -
FFC-AE-V0 [15] 16.313 17.291 0.865 2.669 4.061 3.296 3.373 0.422M 4.064
FFC-AE-V1 [15] 16.899 17.484 0.893 2.868 4.238 3.497 3.573 1.665M 15.672
FFC-UNet [15] 17.851 18.702 0.894 2.962 4.349 3.510 3.682 7.673M 35.552
HiFi++ [1] 17.954 18.936 0.890 2.906 4.271 3.483 3.609 1.706M 2.746
HiFi++(reprod.) 13.400 15.661 0.866 2.595 3.998 3.173 3.305 1.617M 2.096
HiFi w/o spec 12.808 16.155 0.873 2.703 4.094 3.166 3.413 1.356M 2.007
HiFi-2dMRF 13.455 15.812 0.865 2.466 3.865 3.075 3.172 0.677M 2.084
HiFi-Stream 14.067 15.924 0.865 2.701 4.031 3.220 3.372 1.174M 1.895
HiFi-Stream2D 14.762 17.213 0.864 2.652 4.027 3.167 3.345 0.497M 1.973
@@ IV Experimental Setup
Model SI-SDR (dB) $\uparrow$ SDR (dB) $\uparrow$ STOI $\uparrow$ PESQ $\uparrow$ CSIG $\uparrow$ CBAK $\uparrow$ COVL $\uparrow$
Ground Truth - - 1.000 4.640 5 5 5
Input 8.440 8.54 0.790 1.970 3.340 2.820 2.740
DEMUCS(caus.dns)[3] 17.804 19.310 0.881 2.657 3.823 3.231 3.236
FFC-AE-V0 [15] 12.664 13.820 0.789 1.996 3.502 2.807 2.742
FFC-AE-V1 [15] 15.147 15.806 0.860 2.424 3.906 3.214 3.172
FFC-UNet [15] 15.683 16.467 0.861 2.452 3.974 3.225 3.222
HiFi++ [1] 14.669 15.264 0.930 2.365 3.838 3.162 3.103
HiFi++(reprod.) 13.014 15.383 0.841 2.263 3.715 2.917 2.992
HiFi w/o spec 11.839 14.746 0.791 1.932 3.200 2.671 2.559
HiFi-2dMRF 12.771 14.861 0.806 1.931 3.259 2.726 2.587
HiFi-Stream 13.328 15.123 0.845 2.295 3.667 2.860 2.981
HiFi-Stream2D 14.189 16.328 0.847 2.286 3.680 2.925 2.982
Figure 3: Performance metric (COVL) against the computational complexity and number of parameters for compared models (streaming evaluation).
IV-A Data and Evaluation
We used Valentini VCTK [19] dataset for experiments. The dataset is a selection of $30$ speakers from the VoiceBank corpus with $28$ and $2$ speakers set for the training and test respectively. The speech utterances were mixed with noise utterances (exact specification is given in [19]). The original sample rate of the dataset is $48$ kHz, we resampled the audio to $16$ kHz for comparison with other SE solutions. All audios staged for train and validation parts were cut to $3$ seconds while test part is not changed [1].
@@ IV-A Data and Evaluation
For evaluation we used perceptual evaluation metrics representing the quality and intelligibility of the speech: CSIG, CBAK, COVL [1], WB-PESQ [13], STOI [18]. Also, we present objective evaluation metrics: Signal-To-Distortion Ration (SDR, dB)
$SDR(x,\hat{x})=20\log_{10}\frac{\|x\|}{\|x-\hat{x}\|}$ (3)
and its scale-invariant version (SI-SDR, dB) given by [14] to ensure the amplitude invariance. The key aspect of our research is model complexity, so we also demonstrate the complexity metrics: number of parameters and GMACs per second of audio measured with pytorch profiling library thop.
@@ IV-B Training details
All models were trained for 150 epochs with learning rate 2e-4 on the given dataset with batch size 32 using AdamW optimizer (see the repository for the exact details). We used the same loss construction and discriminators as in [1, 7] and focus our main efforts on the generator structure. The generator loss is a composite loss
$\mathcal{L}(\theta)=\mathcal{L}_{GAN}(\theta)+\alpha_{FM}\mathcal{L}_{FM}(\theta)+\alpha_{Mel}\mathcal{L}_{Mel}(\theta)$ (4)
consisting of adversarial, feature-matching and mel reconstruction losses with $\alpha_{FM}=2$ and $\alpha_{Mel}=45$. Discriminators (3 MSD as in [1]) are trained with only optimizing adversarial component
$\mathcal{L}(\phi_{i})=\mathcal{L}_{GAN}(\phi_{i})$ (5)
for $i$-th discriminator. More specific details for reproduction are available in our configuration files.